OpenStudy (alyssa):

how do you find the summation to write a series for (-6,-5,-4,-3,-2)

OpenStudy (anonymous):

\[\sum_{k=-6}^{-2}k\] which is the sum of each real in the set

OpenStudy (anonymous):

You are adding (finding the summation) of the set (-6, -5, -4, -3, -2). Sigma notation adds k, starting at a specific number (-6), and adding each whole in between until it reaches the top number (-2).

OpenStudy (anonymous):

\[a_0=-6, a_n=a_m+1, m = n-1, n = [0, 5]\]